how many 25 cents are there in $17.25? how do you know this? how did you get the answer?
Answer by jim_thompson5910(35256)

x = # of quarters


(# of quarters)*(value of 1 quarter) = total value

x * 0.25 = 17.25

x = 17.25/0.25

x = 69


So you have 69 quarters.
------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------